40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15994     Entry time: Sat Apr 3 00:42:40 2021
Author: gautam 
Type: Update 
Category: LSC 
Subject: PRFPMI locking with half input power 

Summary:

I wanted to put my optomechanical instability hypothesis to the test. So I decided to cut the input power to the IMC by ~half and try locking the PRFPMI. However, this did not improve the stability of the buildup in the arm cavities, while the control was solely on the ALS error signal

Details:

  1. The waveplate I installed for this purpose was rotated until the MC RFPD DCMON channel reported ~half it's nominal value.
  2. I adjusted the IMC servo gains appropriately to compensate. IMC lock was readily realized.
  3. I increased the whitening gains on the POX, POY and REFL165 photodiodes by 6dB, to compensate for the reduced light levels.
    • One day soon, we will have remote power control, and it'd be nice to have this process be automated.
    • Really, we should have de-whitening filters that undo these flat gains in addition to undoing the frequency dependent whitening.
    • I'm not sure the quality of the electronics is good enough though, for the changing electronics offsets to not be a problem.
    • One possibility is that we can normalize some signals by the DC light level at that port, but I still think compensating the changing optical gain as far upstream as possible is best, and the whitening gain is the convenient stage to do this.
  4. Recovered single arm POX/POY locking. 
  5. Then I decided to try and lock the PRFPMI with the reduced input power.

Basically, with some tweaks to loop gains, it worked, see Attachment #1. Note that the lower right axis shows the IMC transmission and is ~7500 cts, vs the nominal ~15,000 cts.

Discussion:

Cutting the input power did not have the effect I hoped it would. Basically, I was hoping to zero the optical CARM offset while the IFO was entirely under ALS control, and have the arm transmission be stable (or at least, stay in the linear regime of REFL11). However, the observation was that the IFO did the usual "buzzing" in and out of the linear regime. Right now, this is not at all a problem - once the IR error signal is blended in, and DC control authority is transferred to that signal, the lock acquisition can proceed just fine. And I guess it is cool that we can lock the IFO at ~half the input power, something to keep in mind when we have the remote controlled waveplate, maybe we always want to lock at the lowest power possible such that optomechanical transients are not a problem. 

I also don't think this test directly disputes my claim that the residual CARM noise when the arm cavities are under purely ALS control is smaller than the CARM linewidth.

What does this mean for my hypothesis? I still think it is valid, maybe the power has to be cut even further for the optomechanics to not be a problem. In Finesse (see Attachment #2), with 0.3 W input power to the back of the PRM, and with best guesses for the 40m optical losses in the PRC and arms, I still see that considerable phase can be eaten up due to the optomechanical resonance around ~100 Hz, which is where the digital CARM loop UGF is. So I guess it isn't entirely unreasonable that the instability didn't go away?


After this work, I undid all the changes I made for the low power lock test. I confirmed that IMC locking, POX/POY locking, and the dither alignment systems all function as expected after I reverted the system.

Attachment 1: PRFPMIlock_1301464998_1301465238.pdf  63 kB  Uploaded Sat Apr 3 02:00:01 2021  | Hide | Hide all
PRFPMIlock_1301464998_1301465238.pdf
Attachment 2: CARMplant.pdf  493 kB  Uploaded Sat Apr 3 02:13:09 2021  | Hide | Hide all
CARMplant.pdf
ELOG V3.1.3-